Fill in the blanks:

Fit Apparel

Boswell, Inc.

Net Sales Revenue

$

99,000

$

(d)

Cost of Goods Sold:

Beginning Merchandise Inventory

(a)

31,000

Purchases and Freight In

47,000

(e)

Cost of Goods Available for Sale

(b)

88,000

Ending Merchandise Inventory

(2,000)

(2,000)

Cost of Goods Sold

60,000

(f)

Gross Profit

39,000

113,000

Selling and Administrative Expenses

(c)

87,000

Operating Income

$

11,000

$

(g)
Answer #1

Fit Apparel

(b) Cost of Goods Sold = Cost of Goods Available for Sale - Ending Merchandise Inventory

60,000 = Cost of Goods Available for Sale - 2,000

Cost of Goods Available for Sale = $62,000

(a) Cost of Goods Available for Sale = Beginning Merchandise Inventory + Purchases and Freight In

62,000 = Beginning Merchandise Inventory + 47,000

Beginning Merchandise Inventory = $15,000

(c) Operating Income = Gross Profit - Selling and Administrative Expenses

11,000 = 39,000 - Selling and Administrative Expenses

Selling and Administrative Expenses = $28,000

Boswell, Inc.

(f) Cost of Goods Sold = Cost of Goods Available for Sale - Ending Merchandise Inventory

= 88,000 - 2,000

= $86,000

(d) Net Sales Revenue = Cost of Goods Sold + Gross Profit

= 86,000 + 113,000

= $199,000

(e) Cost of Goods Available for Sale = Beginning Merchandise Inventory + Purchases and Freight In

88,000 = 31,000 + Purchases and Freight In

Purchases and Freight In = $57,000

(g) Operating Income = Gross Profit - Selling and Administrative Expenses

= 113,000 - 87,000

= $26,000

Please ask if you have any query related to the question. Thank you